References

[1] R.A. Strehlow. Transverse waves in detonations: II. structure and spacing in H_{2}-O_{2}, C_{2}H_{2}-O_{2}, C_{2}H_{4}-O_{2} and CH_{4}-O_{2} systems. AIAA J., 7(3):492– 496, 1969.

[2] R.A. Strehlow. Multi-dimensional detonation wave structure. Astronaut. Acta, 15(5):345–357, 1970.

[3] R.A. Strehlow, R. Liangminas, R.H. Watson, and J.R. Eyman. Transverse wave structure in detonations. In 11th Symp. Int. Combust. Proc., pages 683–692, 1967.